OpenStudy (anonymous):

Find the x-coordinates of any relative extrema and inflection point(s) for the function f(x) = 6x^(1/3) + 3x^(4/3). You must justify your answer using an analysis of f ′(x) and f ′′(x).

OpenStudy (jhannybean):

relative extremas can be determined by finding the first derivative, setting it equal to 0, then finding your critical points and testing where they change signs.

OpenStudy (anonymous):

Well i know the derivative is 2x^-2/3 + 4x^1/3 but i forget how to solve for x here

OpenStudy (jhannybean):

\[\ f(x) = 6x^{\frac{1}{3}} + 3x^{\frac{4}{3}}\]\[\ f'(x) = 6 \cdot \frac{1}{3}x^{-\frac{2}{3}} + 3 \cdot \frac{4}{3} x^{\frac{1}{3}}\]\[\ f'(x) = 2x^{-\frac{2}{3}} + 4x^{\frac{1}{3}} = 0 \] Now you need to factor out the LCM And good job getting the derivative :)

OpenStudy (jhannybean):

Between 2 and 4, 2 is the LCM, and between\(\ -\frac{2}{3} \&\ \frac{1}{3}\) , \(\ -\frac{2}{3}\) is smaller. right?

OpenStudy (anonymous):

so the equation would no be 2x^-2/3(x + 2x^1/2) i presume

OpenStudy (jhannybean):

well, when you multiply \(\ -\frac{2}{3} \cdot \frac{1}{2}\) do you get \(\frac{1}{3}\)? :o

OpenStudy (anonymous):

yea my mistake so 2x^-2/3(x + 2x^-1/2)

OpenStudy (jhannybean):

Yep! and now you set this equal to 0

OpenStudy (jhannybean):

\(\ \implies 2x^{-\frac{2}{3}} = 0 \ \ and\ \ x+ 2x^{-\frac{1}{2}} = 0\)

OpenStudy (anonymous):

This is where i get confused because i forget how to solve these kinds of equations

OpenStudy (jhannybean):

Hmm... one sec.

OpenStudy (jhannybean):

You know.. we were going about the hard way of doing it :(

OpenStudy (jhannybean):

Sorry!! Let's try this again starting with this step. \[\ f'(x) = 6 \cdot \frac{1}{3}x^{-\frac{2}{3}} + 3 \cdot \frac{4}{3} x^{\frac{1}{3}}\] from here, we can rewrite our powers as positive powers as well as reduce our function. \[\ \large f'(x) = \frac{2}{x^{\frac{2}{3}}} + 4x^{\frac{1}{3}}\]

OpenStudy (jhannybean):

Now we find a common denominator, \(\ x^{\frac{2}{3}}\)

OpenStudy (jhannybean):

So we'll end up with \[\ \large f'(x) = \frac{4x + 2}{x^{\frac{2}{3}}}\] :) From here...

OpenStudy (anonymous):

Ohhhh!!! this makes much more sense so the critical points are know x = 0, .5

OpenStudy (jhannybean):

We set \[\ \large x^{\frac{2}{3}} = 0 \implies x=0\] \[\ 4x+2 = 0 \implies x =\frac{1}{2} \]

OpenStudy (jhannybean):

Yesss you are right :)

OpenStudy (jhannybean):

and so those are your critical points of the relative extrema, and for the inflection points, we have to find the second derivative

OpenStudy (jhannybean):

we can use the product rule! All we have to do is rewrite our derivative in a multiplication form :) \[\ \large f'(x) = (4x+2)(x^{\frac{2}{3}})\] Then we just apply the product rue to this "\(\ f''g + g'f\)"

OpenStudy (anonymous):

wouldn't it be the quotient rule?

OpenStudy (jhannybean):

I meant \[\ f'(x) = (4x+2)(x^{-\frac{2}{3}})\]

OpenStudy (anonymous):

I see.

OpenStudy (jhannybean):

You can use the quotient rule if you'd like to leave your derivative as it is, then \(\ f = 4x+2 \ \ and \ \ g = x^{\frac{2}{3}}\) and you just apply the quotient rule, that is \(\ \frac{f'g - g'f}{g^2}\)

OpenStudy (jhannybean):

But I prefer the product rule since I don't like dealing with fractions :P

OpenStudy (jhannybean):

Oh, i typoed the product rule too D: Grr. It's f'g + g'f **

OpenStudy (anonymous):

so f''(x) = 4x^-2/3 + -2/3x^-5/3(4x + 2)

OpenStudy (jhannybean):

Not the second derivative of f x_x

OpenStudy (jhannybean):

\[\ \large f'(x) = (4x+2)(x^{-\frac{2}{3}})\]\[\ f'g + g'f\]\[\ 4(x^{-\frac{2}{3}}) + \left(-\frac{2}{3}\right)x^{-\frac{5}{3}} \cdot (4x+2)\]

OpenStudy (jhannybean):

Solving this i'm realizing there is no easy way of going about finding your inflection points :(

OpenStudy (anonymous):

I think i can figure it out from here and you have my sincerest thanks for helping me

OpenStudy (jhannybean):

And let's try the quotient rule with this.

OpenStudy (jhannybean):

If we have.... \[\ \large f'(x) = \frac{4x + 2}{x^{\frac{2}{3}}}\] Then the quotient rule will give us \[\frac{4(x^{\frac{2}{3}})-\left(\frac{2}{3}\right)x^{-\frac{1}{3}}(4x+2)}{\left(x^{\frac{2}{3}}\right)^2}\]

OpenStudy (jhannybean):

It will reduce down to \[\large -\frac{4}{3x^{\frac{5}{3}}} + 4\left( \frac{1}{3x^{\frac{2}{3}}}\right)\] which will further reduce to \[f''(x) =\frac{4(x-1)}{3x^{\frac{5}{3}}}\]

OpenStudy (jhannybean):

From here you just follow the same steps in finding your critical points, as we did earlier with the first derivative, and you'll find your inflection points

OpenStudy (jhannybean):

Good luck!